### v1

# The Ehrhart series of alcoved polytopes

Abstract.
Alcoved polytopes are convex polytopes which are the closure of a union of alcoves in an affine Coxeter arrangement. They are rational polytopes, and therefore have Ehrhart quasipolynomials. We describe a method for computing the generating function of the Ehrhart quasipolynomial, or Ehrhart series, of any alcoved polytope via a particular shelling order of its alcoves. We also conjecture a connection between Early’s decorated ordered set partitions and this shelling order for the hypersimplex $\Delta_{2,n}$.
1. Introduction
Let $\Phi\subset V$ be an irreducible crystallographic root system and $W$ be the corresponding Weyl group. Associated to $\Phi$ is an infinite hyperplane arrangement known as the affine Coxeter arrangement. The connected components of this hyperplane arrangement are simplices called alcoves. Lam and Postnikov defined a proper alcoved polytope to be the closure of a union of alcoves.
In the special situation of the root system $\Phi=A_{n}$, examples of alcoved polytopes include hypersimplices and positroid polytopes. For any root system, Fomin and Zelevinsky’s generalized associahedra are examples of polytopes which can be realized as alcoved polytopes [fz03, chapoton].
Alcoved polytopes are rational polytopes, i.e., their vertices have rational coordinates. Let $P\subset\mathbb{R}^{n}$ be a rational convex polytope. Ehrhart [ehrhart] showed that the number of lattice points in the $t$-dilate of $P$ is a quasipolynomial in $t$. A quasipolynomial with period $d$ is a function $p:\mathbb{Z}\to\mathbb{Z}$ such that there exist periodic functions $p_{i}:\mathbb{Z}\to\mathbb{Z}$ with period $d$ so $p(z)=\sum_{i=0}^{n}p_{i}(z)z^{i}$. We call $E(P,t)=\#(t\cdot P)\cap\mathbb{Z}^{n}$ the Ehrhart quasipolynomial of $P$. The generating series $\sum_{t=0}^{\infty}E(P,t)z^{t}$ is a rational function in $z$. While the Ehrhart theory of lattice polytopes has been widely studied, the Ehrhart theory of rational polytopes is a younger subject [beck22, BBKV13, linke11].
Ehrhart theory extends naturally to polytopes with some facets removed. The Ehrhart series are additive, in the sense that if two half-open polytopes are disjoint, then the Ehrhart series of their union is equal to the sum of their Ehrhart series.
Our main result stated imprecisely is the following:
Theorem.
Fix an irreducible crystallographic root system $\Phi\subset V$, where $\dim(V)=n$. Let $P$ be an alcoved polytope and let $\Gamma_{P}=(V,E)$ be the dual graph to the alcove triangulation of $P$. Pick some $v_{0}\in V$ and orient the edges of $\Gamma_{P}$ so that for all $\{u,w\}\in E$, $u\to w$ if and only if $u$ appears before $v$ in the breadth first search algorithm starting at $v_{0}$. There exists a weighting of the edges $E$ and parameters $\ell_{1},\cdots,\ell_{n}$ depending only on $\Phi$ such that the Ehrhart series of $P$ is equal to
$\operatorname{Ehr}(P,z)=\frac{\sum_{w\in V}z^{\operatorname{wt}(w)}}{\prod_{i=0}^{n}(1-z^{\ell_{i}})}$
where $\operatorname{wt}(w)=\sum_{u\to w}\operatorname{wt}((u,w))$ is the sum of the weights of the ingoing edges to $w$.
We will prove our main result using the additivity of Ehrhart series. We will decompose each alcoved polytopes into disjoint union of half-open alcoves, and then add up all their Ehrhart series.
(0,0)(1,0)($\frac{1}{2},\frac{1}{2}$)
[…]
Figure 1. On the left, we have the fundamental alcove of type $B_{2}$. On the right, we have an alcoved polytope of type $B_{2}$ and the graph of its alcoved triangulation. The Ehrhart series of the square is $\frac{1+z+z^{2}}{(1-z)^{2}(1-z^{2})}$.
$(0,0,0)$$(\frac{1}{2},0,-\frac{1}{2})$$(\frac{2}{3},-\frac{1}{3},-\frac{1}{3})$
[…]
Figure 2. On the left, we have the fundamental alcove of type $G_{2}$. On the right, we have an alcoved polytope of type $G_{2}$. The Ehrhart series of the trapezoid is $\frac{1+z^{2}+z^{3}}{(1-z)(1-z^{2})(1-z^{3})}$.
After proving this result, we conjecture a relationship between this formula and another formula for the $h^{*}$-polynomial of the second hypersimplex $\Delta_{2,n}=\{(x_{1},\dots,x_{n})\in[0,1]^{n}~{}|~{}\sum_{i=1}^{n}x_{i}=2\}$ which is given in terms of combinatorial objects called decorated ordered set partitions (DOSPs). The only known proofs of the latter formula rely on algebraic manipulation and enumerative combinatorics; our conjecture is aimed towards providing a more geometric reason for why DOSPs appear in the $h^{*}$-polynomial of the hypersimplex.
1.1. Organization
In Section 2, we cover relevant background materials, including root systems, alcoved polytopes, and Coxeter groups. In Section 3, we review some general Ehrhart theory of rational polytopes and prove a few lemmas that will be used to show our main result. In Section 4, we review Coxeter complexes and the notion of a convex subset of a Coxeter group. In this section, we also describe a shelling order of the Coxeter complex given by [bjorner] and prove a corollary about shelling subcomplexes of the Coxeter complex (Corollary 4.6) which will be used to show our main result. We precisely state and prove our main result (Theorem 5.4) in Section 5. In Section 6, we recall a formula for the $h^{*}$-polynomial of the hypersimplex $\Delta_{k,n}$ in terms of hypersimplicial decorated ordered set partitions [kim2020combinatorial] and conjecture a connection with the formula yielded by our main theorem in the case $k=2$.
Acknowledgements
We thank Alex Postnikov for suggesting this topic. We thank Lauren Williams for the helpful discussions and comments on the manuscript. We thank Nick Early for helpful discussions.
2. Preliminaries
In this section, we recall the relevant background for root systems which will be used to define alcoved polytopes. We recall notations from Coxeter groups which will be used to describe our shellings of the Coxeter complexes related to the alcove triangulation of an alcoved polytope. We follow the conventions in [humphreys] and [alcove2].
2.1. Root systems
Let $V$ be a real Euclidean space of rank $n$ with nondegenerate symmetric inner product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ). Let $\Phi\subset V$ be an irreducible crystallographic root system with a choise of basis of simple roots $\alpha_{1},\dots,\alpha_{n}$. Let $\Phi^{+}\subset\Phi$ be the corresponding set of positive roots. The coweight lattice $\Lambda^{\vee}$ is the integer lattice defined by $\Lambda^{\vee}=\Lambda^{\vee}(\Phi)=\{\lambda\in V\mid(\lambda,\alpha)\in\mathbb{Z},\text{ for all }\alpha\in\Phi\}$. Let $\omega_{1},\dots,\omega_{n}\subset V$ be the basis dual to the basis of simple roots, i.e., $(\omega_{i},\alpha_{j})=\delta_{ij}$. The $\omega_{i}$ are called the fundamental coweights. They generate the coweight lattice $\Lambda^{\vee}$.
Let $\rho=\omega_{1}+\cdots+\omega_{n}$. The height of a root $\alpha$ is the number $(\rho,\alpha)$ of simple roots that add up to $\alpha$. Since we assumed that $\Phi$ is irreducible, there exists a unique highest root $\theta\in\Phi^{+}$ of maximal possible height. For convenience we set $\alpha_{0}=-\theta$. Let $a_{0}=1$ and $a_{1},\dots,a_{n}$ be the positivie integers given by $a_{i}=(\omega_{i},\theta)$. The dual Coxeter number is defined as $h^{\vee}=(\rho,\theta)+1=a_{0}+a_{1}+\cdots+a_{n}$.
2.2. Alcoved polytopes
Let $\Phi\subset V$ be a crystallographic root system. The set of affine hyperplanes of the form $H_{\alpha,k}=\{\lambda\in V\mid(\lambda,\alpha)=k\}$, where $\alpha\in\Phi^{+},k\in\mathbb{Z}$, divides $V$ into open alcoves:
Definition 2.1.
An (open) alcove is the set
$A=\{\lambda\in V\mid m_{\alpha}<(\lambda,\alpha)<m_{\alpha}+1,\text{ for }\alpha\in\Phi^{+}\}$
where $m_{\alpha}$ is a collection of integers associated with the alcove $A$. A closed alcove is the closure of an alcove.
Definition 2.2.
The fundamental alcove is the simplex given by
$A_{\circ}$ =$=\{\lambda\in V\mid 0<(\lambda,\alpha)<1,\text{ for }\alpha\in\Phi^{+}\}$
$=\text{Convex Hull of the points }0,\omega_{1}/a_{1},\dots,\omega_{r}/a_{r}.$
The affine Weyl group $W_{\mathrm{aff}}$ associated with the root system $\Phi$ is generated by the reflections $s_{\alpha,k}:V\to V,\alpha\in\Phi,k\in\mathbb{Z}$, with respect to the affine hyperplanes $H_{\alpha,k}$, and acts simply transitively on the collection of all alcoves [humphreys]. In particular, the closure of each alcove has the same Ehrhart series, and we can use alcoves as the building blocks for a family of polytopes with natural triangulations in which all top-dimensional simplices have equal volume:
Definition 2.3.
An alcoved polytope is a polytope which is the union of some collection of faces of alcoves. A proper alcoved polytope is a polytope which is a union of alcoves, equivalently a top-dimensional alcoved polytope.
In other words, every alcoved polytope is of the form
$P=\{\lambda\in V\mid k_{\alpha}\leq(\lambda,\alpha)\leq K_{\alpha},\text{ for }\alpha\in\Phi^{+}\},$
where $k_{\alpha},K_{\alpha}$ are two collections of integers indexd by the positive roots $\alpha\in\Phi^{+}$.
Definition 2.4.
Let $P$ be an alcoved polytope. We associate a graph $\Gamma_{P}=(V,E)$ with labeled edges to the alcoved triangulation of $P$. We will abuse notation and also use $\Gamma_{P}$ to denote the simplicial complex of the alcove triangulation of $P$. The vertex set $V$ consists of closed alcoves in $P$, and the edge set $E$ consists of $(A,A^{\prime})$ if $A$ and $A^{\prime}$ share a common facet.
Let $(A,A^{\prime})$ be an edge of $G$, and let $F=A\cap A^{\prime}$ be the facet it represents. Then $F$ can be transformed to a facet $F_{\circ}$ of the fundamental alcove $A_{\circ}$ under the action of the affine Weyl group. Let $\omega_{i}/a_{i}$ be the vertex of $A_{\circ}$ that does not belong to $F_{\circ}$. Then $(A,A^{\prime})$ has weight $\ell_{i}$, denoted $\operatorname{wt}((A,A^{\prime}))=\ell_{i}$, in which $\ell_{i}$ is the least common multiple of the denominators in $\omega_{i}/a_{i}\in\mathbb{Q}^{n}$ for $i=1,\dots,n$ and $\ell_{0}=1$.
2.3. Coxeter Groups
For a root system $\Phi$, the affine Weyl group $W_{\mathrm{aff}}$ is an example of a Coxeter group:
Definition 2.5.
Let $S$ be a set. A matrix $m:S\times S\to\{1,2,\dots,\infty\}$ is called a Coxeter matrix if it satisfies
$m(s,s^{\prime})$ $=m(s^{\prime},s)$
$m(s,s^{\prime})=1$ $\iff s=s^{\prime}.$
A Coxeter matrix $m$ determines a group $W$ with the presentation
$\langle s\in S\mid(ss^{\prime})^{m(s,s^{\prime})}=e\text{ for all }m(s,s^{\prime})\neq\infty\rangle,$
where $e$ is the identity element. The pair $(W,S)$ is called a Coxeter system and the group $W$ is the Coxeter group.
Definition 2.6.
Let $(W,S)$ be a Coxeter system. Each element $w\in W$ can be written as a product of generators $w=s_{1}s_{2}\cdots s_{k}$ for $s_{i}\in S$. If $k$ is minimal among all such expressions for $w$, then $k$ is called the length of $w$ (written $\ell(w)=k$) and the word $s_{1}s_{2}\cdots s_{k}$ is called a reduced word for $w$.
Definition 2.7.
Let $(W,S)$ be a Coxeter system, and let $u,v\in W$. Define the right weak order as the partial order $\leq$ on $W$ such that $u\leq w$ if and only if $w=us_{1}s_{2}\cdots s_{k}$ for some $s_{i}\in S$ such that $\ell(us_{1}s_{2}\cdots s_{i})=\ell(u)+i$ for all $0\leq i\leq k$.
Definition 2.8.
Let $W$ be a Coxeter group. For $w\in W$, define the descent set $\mathcal{D}(w)=\{s\in S\mid ws<w\}$. For $J\subseteq S$, let $W_{J}$ be the subgroup of $W$ generated by the set $J$, and let $W^{J}=\{w\in W\mid ws>w\text{ for all }s\in J\}$ be a system of distinct coset representatives modulo $W_{J}$.
3. Ehrhart series of half-open rational simplices
A key ingredient of the proof of our main theorem is to write an alcoved polytope $P$ as a disjoint union of half-open simplices, or simplices with some facets removed. Since the number of lattice points in $P_{1}\sqcup P_{2}$ is the sum of lattice points in $P_{1}$ and $P_{2}$, it suffices to look at the Ehrhart series of these half-open simplices. If we remove no facets from an alcove, we may use the following theorem to write its Ehrhart series:
Theorem 3.1 (Theorem 1.3, [stanley80]).
Suppose $S$ is a $k$-simplex in $\mathbb{R}^{m}$ with rational vertices $\beta_{0},\dots,\beta_{k}$. Let $\ell_{i}$ be the least positive integer $t$ for which $t\beta_{i}\in\mathbb{Z}^{m}$. Consider the $(k+1)\times(m+1)$-matrix whose rows are the vectors $(\ell_{i}\beta_{i},\ell_{i})$. If the greatest common divisor of all $(k+1)\times(k+1)$ minors of the matrix is equal to 1, then the Ehrhart series of the simplex $S$ is equal to
$\operatorname{Ehr}(S,z)=\frac{1}{\prod_{i=0}^{k}(1-z^{\ell_{i}})}.$
The above theorem serves as a base case in the proof of the following more general formula for half-open simplices:
Lemma 3.2.
Suppose $S$ is a $k$-simplex in $\mathbb{R}^{m}$ with rational vertices $\beta_{0},\dots,\beta_{k}$. Let $\ell_{i}$ be the least positive integer $t$ for which $t\beta_{i}\in\mathbb{Z}^{m}$. Let $F_{i}$ be the facet of $S$ that does not contain the vertex $\beta_{i}$. Let $\mathcal{F}\subseteq\{0,\dots,k\}$ label a subset of facets of $S$. Then the Ehrhart series of $S$ with facets $\{F_{i}\mid i\in\mathcal{F}\}$ removed is
$\operatorname{Ehr}(S\setminus\bigcup_{i\in\mathcal{F}}F_{i},z)=\frac{\prod_{i\in\mathcal{F}}z^{\ell_{i}}}{\prod_{i=0}^{k}(1-z^{\ell_{i}})}.$
Proof.
Without loss of generality assume =$\mathcal{F}=\{0,\ldots,m\}$. We induct on $m$. First, if we remove no facets, we are done by Theorem 3.1. By inclusion-exclusion and our inductive hypothesis,
$\operatorname{Ehr}\left(S\setminus\bigcup_{i\in\mathcal{F}}F_{i},z\right)$ $=\operatorname{Ehr}(S,z)+\sum_{\mathcal{G}\subseteq\mathcal{F}}(-1)^{|\mathcal{G}|}\operatorname{Ehr}\left(\bigcap_{F\in\mathcal{G}}F,z\right)$
$=\frac{1}{\prod_{i=0}^{k}(1-z^{\ell_{i}})}+\sum_{\mathcal{G}\subseteq\mathcal{F}}\frac{(-1)^{|\mathcal{G}|}}{\prod_{i\in V(\mathcal{G})}(1-z^{\ell_{i}})}$
where $V(\mathcal{G})$ is the set of indices of vertices $\bigcap_{F\in\mathcal{G}}F$. We have $V(\mathcal{G})=\{0,\ldots,k\}\setminus\mathcal{G}$. Clearing denominators,
$\operatorname{Ehr}\left(S\setminus\bigcup_{i\in\mathcal{F}}F_{i},z\right)$ $=\frac{1}{\prod_{i=0}^{k}(1-z^{\ell_{i}})}\left[\sum_{\mathcal{G}\subseteq\mathcal{F}}(-1)^{|\mathcal{G}|}\prod_{i\in\mathcal{G}}(1-z^{\ell_{i}})\right].$
Using inclusion-exclusion again, the last sum is equal to $\prod_{i\in\mathcal{F}}z^{\ell_{i}}$.
4. A shelling order
The goal of this section is to build up to a concrete decomposition of any alcoved polytope into half-open alcoves. This is done via a shelling order. In particular, we relate the alcove structure to a simplicial complex called the coxeter complex and use existing results about the latter to define a shelling order.
We begin by recalling the notion of Coxeter complexes, following [bjornerbrenti].
Definition 4.1.
An abstract simplicial complex $\Delta$ on the vertex set $V$ is a collection $\Delta$ of finite subsets of $V$, called faces, such that $x\in V$ implies $\{x\}\in\Delta$ and $F\subseteq F^{\prime}\in\Delta$ implies $F\in\Delta$. The dimension of a face is $\dim F=|F|-1$, and $\dim\Delta=\sup_{F\in\Delta}\dim F$. A complex is pure $d$-dimensional if every face is contained in some $d$-dimensional face. In this case, we denote the set of $d$-dimensional faces (called facets) by $\mathscr{F}(\Delta)$. Two facets $A,A^{\prime}$ are adjacent if $\dim(A\cap A^{\prime})=d-1$. If $F\in\Delta$, let $\bar{F}$ be the simplex $\{E\mid E\subseteq F\}$.
Definition 4.2.
Let $(W,S)$ be a Coxeter system with $|S|<\infty$. We write $(s):=S\setminus\{s\}$, for $s\in S$, and $V:=\bigcup_{s\in S}W/W_{(s)}$ for the collection of all left cosets of all maximal parabolic subgroups. The Coxeter complex $\Delta(W,S)$ is by definition the pure $(|S|-1)$-dimensional simplicial complex on the vertex set $V$ with facets $C_{w}:=\{wW_{(s)}\mid s\in S\}$, for $w\in W$. For a face $F\in\Delta(W,S)$, define its type $\tau(F)=\{s\in S\mid F\cap W/W_{(s)}\neq\emptyset\}$.
Let $(W,S)$ be a Coxeter system and $u,v\in W$. A path from $u$ to $v$ is a sequence $u=w_{0}\to w_{2}\to\cdots\to w_{s}=v$ such that $w_{i+1}=w_{i}s$ for some simple reflection $s\in S$. A subset $K\subseteq W$ is called convex if for every $u,v\in K$ we have that any shortest path from $u$ to $v$ lies in $K$. This mirrors the notion of convexity for unions of closed alcoves:
Proposition 4.3 (Prop. 3.5, [alcove2]).
Let $P$ be a bounded subset which is a union of closed alcoves. Then $P$ is a convex polytope if and only if one of the following conditions hold:
  1. (1)
    For any two alcoves $A,B\subset P$, any shortest path from $A=A_{0}\to A_{1}\to A_{2}\to\cdots\to A_{s}=B$ lies in $P$. Here $A_{i}$ are alcoves and $A^{\prime}\to A^{\prime\prime}$ means that the closure of the two alcoves $A^{\prime}$ and $A^{\prime\prime}$ share a facet.
  2. (2)
    The subset $W_{P}=\{w\in W_{\mathrm{aff}}\mid w(A_{\circ})\subset P\}$ of the affine Weyl group is a convex subset.
Let $(W,S)$ be a Coxeter system and let $\mathscr{F}(\Delta(W,S))$ denote the set of all facets of $\Delta(W,S)$. By [bourbaki, pp. 40-44] and [tits, Chap. 2], there is a bijection between $W$ and $\mathscr{F}(\Delta(W,S))$ given by $w\mapsto C_{w}$. Two facets $C_{w}$ and $C_{w^{\prime}}$ are adjacent if and only if $w^{\prime}=ws$ for some $s\in S$. By [humphreys], there is a bijection between $\mathscr{F}(\Delta(W,S))$ and the set of alcoves in the affine Coxeter arrangement. This bijection maps $C_{e}$ to the fundamental alcove $A_{\circ}$ (see Definition 2.2), and maps $C_{w}$ to the alcove $A$ such that there is a shortest path $A_{\circ}\overset{s_{1}}{\to}A_{1}\overset{s_{2}}{\to}\cdots\overset{s_{k}}{\to}A$ where $s_{1}s_{2}\cdots s_{k}$ is a reduced word for $w$. The group $W$ acts on $\Delta(W,S)$ by left translation $w:vW_{(s)}\mapsto wvW_{(s)}$, and this action is type-preserving, i.e., $\tau(w(F))=\tau(F)$ for all faces $F\in\Delta(W,S)$.
The key property of Coxeter complexes of interest in this paper is that they have a shelling order:
Definition 4.4.
Let $\Delta$ be a pure $d$-dimensional complex of at most countable cardinality. A shelling of $\Delta$ is a linear order $A_{1},A_{2},A_{3},\dots$ on the set of facets of $\Delta$ such that $\bar{A_{k}}\cap\Delta_{k-1}$ is pure $(d-1)$-dimensional for $k=2,3,\dots$, where $\Delta_{k-1}=\bar{A}_{1}\cup\cdots\cup\bar{A}_{k-1}$.
Given a shelling, define the restriction of a facet $A_{k}$ by
$\mathscr{R}(A_{k})=\{x\in A_{k}\mid A_{k}-\{x\}\in\Delta_{k-1}\}.$
Theorem 4.5 (Theorem 2.1, [bjorner]).
Let $(W,S)$ be a Coxeter system, $|S|<\infty$. Then any linear extension of the weak ordering of $W$ assigns a shelling order to the facets of $\Delta(W,S)$.
In particular, Björner showed that the restriction of this shelling is
$\mathscr{R}(C_{w})=\{wW_{(s)}\mid s\in\mathcal{D}(w)\}.$
Corollary 4.6.
Let $(W,S)$ be a Coxeter system, $|S|<\infty$. Let $\Gamma_{P}$ be the subcomplex of $\Delta(W,S)$ induced by a convex subset $P$ of the Coxeter group $W$ that contains the identity element $e$. Any linear extension of the weak order is a shelling of $\Gamma_{P}$.
Proof.
Since $P$ contains the identity and is convex, if $w\in P$, then all the shortest paths from $e$ to $w$ are contained in $P$. That is, all the reduced words of $w$ are contained in $P$. Let $C_{1},C_{2},\dots$ denote a shelling of $\Delta(W,S)$, and let $C_{a_{1}},C_{a_{2}},\dots$ denote the subsequence of $C_{1},C_{2},\dots$ consisting of facets that are in $\Gamma_{P}$. For $k\geq 2$, let $\Delta_{k-1}=\bar{C}_{1}\cup\cdots\cup\bar{C}_{k-1}$ and let $\Delta^{\prime}_{k-1}=\bar{C}_{a_{1}}\cup\cdots\cup\bar{C}_{a_{k-1}}$. Then by definition, $\Delta^{\prime}_{k-1}\subseteq\Delta_{k-1}$, so $C_{a_{k}}\cap\Delta^{\prime}_{k-1}\subseteq C_{a_{k}}\cap\Delta_{{a_{k}}-1}$. Suppose $C_{a_{k}}=C_{w}$ for some $w\in W$. Since all the reduced words of $w$ are contained in $P$, for each $s\in\mathcal{D}(w)$, we have $C_{w}-\{wW_{(s)}\}\in\Delta^{\prime}_{k-1}$, so $C_{a_{k}}\cap\Delta^{\prime}_{k-1}=C_{a_{k}}\cap\Delta_{{a_{k}}-1}$ is pure $(|S|-1)$-dimensional, concluding the proof. ∎
$(1,0,0)$$(1,1,0)$$(\frac{1}{2},\frac{1}{2},0)$$(1,1,1)$(12,12,$(\frac{1}{2},\frac{1}{2},\frac{1}{2})$$(\frac{3}{2},\frac{1}{2},0)$$(\frac{3}{2},\frac{1}{2},\frac{1}{2})$
[…]
Figure 3. The generalized hypersimplex for $\Phi=B_{3}$ and $k=2$. The Ehrhart series of $\Delta^{B_{3}}_{2}$ is $\operatorname{Ehr}(\Delta^{B_{3}}_{2},z)=\frac{1+z+3z^{2}+z^{3}}{(1-z)^{2}(1-z^{2})^{2}}$
5. Proof of the main theorem
We can now translate the shelling order of subcomplexes of the Coxeter complex in Corollary Corollary 4.6 into a shelling order of the alcoves of an alcoved polytope $P$, and use this shelling order to compute the Ehrhart series of $P$.
Definition 5.1.
Let $\Gamma=(V,E)$ be an undirected graph, and let $v_{0}\in V$ be an arbitrary vertex of $\Gamma$. Define the breadth-first search order of $\Gamma$ with root $v_{0}$ as the partial order $(\mathcal{P}_{v_{0},\Gamma},\prec)$ on $V$ such that for two distinct vertices $u,v\in V$, $u\prec v$ if and only if there is a shortest path from $v_{0}$ to $v$ passing through $u$.
The following is a corollary of Corollary 4.6.
Corollary 5.2.
Let $P$ be an alcoved polytope and let $\Gamma_{P}=(V,E)$ be the graph of the alcoved triangulation of $P$. For any $v_{0}\in V$, any linear extension of the partial order $(\mathcal{P}_{v_{0},\Gamma_{P}},\prec)$ is a shelling order of the alcove triangulation of $P$.
Proof.
Let $W=W_{\mathrm{aff}}$ be the affine Weyl group that acts on the alcoves in $P$. Let $W_{P}=\{w\in W\mid w(A_{\circ})\subset\}$, which is convex by Proposition 4.3. Let $v_{0}$ be the alcove $C_{w}$ for some $w\in W$. Then $w^{-1}(W_{P})$ is a convex subset of $W$ that contains the identity, and $w^{-1}(\mathcal{P}_{v_{0},\Gamma_{P}},\prec)$ is the weak order on the facets of $w^{-1}(\Gamma_{P})$. By Corollary 5.2, any linear extension of $w^{-1}(\mathcal{P}_{v_{0},\Gamma_{P}},\prec)$ is a a shelling order of $w^{-1}(\Gamma_{P})$. Since $W$ acts transitively on the set of all alcoves, we conclude the proof. ∎
Lemma 5.3.
Let $P$ be an alcoved polytope. Let $\Gamma_{P}=(V,E)$ be the graph of the alcoved triangulation of $P$. Fix some $v_{0}\in V$ and let $(\mathcal{P}_{v_{0},\Gamma_{P}},\prec)$ be the breadth-first search order on $\Gamma_{P}$ with root $v_{0}$. For each alcove $A$ in $P$, let $\mathscr{I}_{A}=\{\bar{A}\cap\bar{A^{\prime}}\mid A^{\prime}\prec\cdot~{}A\}$ be a subset of facets of the closure $A$. Then the set of half-open alcoves $A^{\circ}=\bar{A}\setminus(\cup\mathscr{I}_{A})$, are mutually disjoint, and their union is equal to $P$, i.e.,
$P=\bigsqcup_{A\in V}A^{\circ}.$
Proof.
We prove this by induction. Let $A\in V$ be an alcove in $P$. Let $\mathcal{A}=\{A^{\prime}\in V\mid A^{\prime}\prec A\}$ be the set of alcoves that comes before $A$ in the topological sort of $\Gamma_{P}$. We show that
  1. (1)
    $A^{\circ}\cap\bigcup_{A^{\prime}\in\mathcal{A}}A^{\prime\circ}=\emptyset$;
  2. (2)
    $A^{\circ}\cup\bigcup_{A^{\prime}\in\mathcal{A}}A^{\prime\circ}=\bar{A}\cup(\bigcup_{A^{\prime}\in\mathcal{A}}\bar{A^{\prime}})$.
To show (1), assume there exists $A^{\prime}\in\mathcal{A}$ such that $A^{\circ}\cap A^{\prime\circ}\neq\emptyset$. Then $\bar{A}\cap\bar{A}^{\prime}\neq\emptyset$ and it must be contained in a facet of $\bar{A}$ in $\bar{A}\cap\bigcup_{A^{\prime}\in\mathcal{A}}\bar{A}^{\prime}$ because $\prec$ is a shelling. This facet is equal to $\bar{A}\cap\bar{A}^{\prime}$ for $A^{\prime}\prec\cdot~{}A$, and therefore will be excluded in $A^{\circ}$.
To show (2), we first use the induction hypothesis that $\bigcup_{A^{\prime}\in\mathcal{A}}A^{\prime\circ}=\bigcup_{A^{\prime}\in\mathcal{A}}\bar{A}^{\prime}$. Then, since $\cup\mathscr{I}_{A}\subset\bigcup_{A^{\prime}\in\mathcal{A}}\bar{A}^{\prime}$, we have $A^{\circ}\cup(\bigcup_{A^{\prime}\in\mathcal{A}}\bar{A}^{\prime})=\bar{A}\cup(\bigcup_{A^{\prime}\in\mathcal{A}}\bar{A}^{\prime})$. ∎
We now have all the ingredients we need to compute the Ehrhart polynomial of an arbitrary alcoved polytope:
Theorem 5.4.
Let $P$ be an alcoved polytope and let $\Gamma_{P}=(V,E)$ be the edge-weighted graph of its alcoved triangulation (see Definition 2.4 for details). Given any $v_{0}\in V$, let $\mathcal{P}_{v_{0},\Gamma_{P}}$ be the breadth-first search order of $\Gamma_{P}$ with root $v_{0}$ (see Definition 5.1). The Ehrhart series of $P$ is equal to
$\operatorname{Ehr}(P,z)=\frac{\sum_{v\in V}z^{\operatorname{wt}(v)}}{\prod_{i=0}^{n}(1-z^{\ell_{i}})}$
where $\operatorname{wt}(v)=\sum_{u\prec\cdot~{}v}\operatorname{wt}((u,v))$ is the sum of the weights of the edges between $v$ and the elements it covers.
Proof of Theorem 5.4.
By Lemma 5.3 and Lemma 3.2 and the additivity of Ehrhart series, we conclude the proof. ∎
Example 5.5.
[…]
6. The hypersimplex $\Delta_{2,n}$
We now speculate on a relationship between our shelling order formula and a combinatorial formula for a well-studied polytope.
The hypersimplex $\Delta_{k,n}$ is the subset of $[0,1]^{n}\subset\mathbb{R}^{n}$ consisting of points $\{(x_{1},\dots,x_{n})\mid x_{1}+\cdots+x_{n}=k\}$. Under the linear transformation
$y_{i}=x_{1}+\cdots+x_{i},$
the hypersimplex $\Delta_{k,n}$ can be realized as an alcoved polytope defined by $0\leq y_{i}-y_{i-1}\leq 1$ and $y_{n}=k$ for all $i=1,\dots,n$ with the convention $y_{0}=0$.
The coefficients of the $h^{*}$-polynomial form the $h^{*}$-vector. The $h^{*}$-vector of the hypersimplex $\Delta_{k,n}$ was proved to be enumerated by hypersimplicial decorated ordered set partitions.
Definition 6.1 ([early2017conjectures]).
A decorated ordered set partition $((S_{1})_{s_{1}},\dots,(S_{p})_{s_{p}})$ of type $(k,n)$ consists of an ordered set partition $(S_{1},\dots,S_{p})$ of $[n]$ and an $p$-tuple of integers $(s_{1},\dots,s_{p})\in\mathbb{Z}^{p}$ such that $\sum_{i=1}^{p}s_{i}=k$ and $s_{i}\geq 1$. We call each $S_{i}$ a block and we place them on a circle in the clockwise fashion then think of $s_{i}$ as the clockwise distance between adjacent block $S_{i}$ and $S_{i+1}$.
We regard them up to cyclic rotation of blocks, so $((S_{1})_{s_{1}},(S_{2})_{s_{2}},\dots,(S_{p})_{s_{p}})$ is the same as $((S_{2})_{s_{2}},\dots,(S_{p})_{s_{p}},(S_{1})_{s_{1}})$.
A decorated ordered set partition is hypersimplicial if $11\leq s_{i}\leq|S_{i}|-$ 1 for all $i$. The winding vector of a decorated ordered set partition is an $n$-tuple of integers $(l_{1},\dots,l_{n})$ such that $l_{i}$ is the distance of the path starting from the block containing $i$ to the block containing $(i+1)$ moving clockwise. If $i$ and $(i+1)$ are in the same block then $l_{i}=0$. If $l_{1}+\cdots+l_{n}=kd$, then we define the winding number to be $d$.
Theorem 6.2 ([Kimh*]).
Let $h^{*}(\Delta_{k,n},z)=h_{0}^{*}(\Delta_{k,n})+h_{1}^{*}(\Delta_{k,n})z+\cdots+h_{n-1}^{*}(\Delta_{k,n})z^{n-1}$ be the $h^{*}$-polynomial of the hypersimplex $\Delta_{k,n}$. The number of hypersimplicial decorated ordered set partitions of type $(k,n)$ and winding number $d$ is $h_{d}^{*}(\Delta_{k,n})$.
In this section, we conjecture the relation between our formula in Theorem 5.4 for the hypersimplex $\Delta_{2,n}$ to the formula of Early and Kim given by the hypersimplicial ordered set partitions.
Definition 6.3.
If a hypersimplicial decorated ordered set partition of type $(2,n)$ with nonzero winding number is of the form $((A)_{1},(B)_{1})$ for nonempty disjoint subsets $A,B\subseteq[n]$ with $A\cup B=[n]$, then we simply denote it by the set of sets $\{A,B\}$. This notation also emphasizes the fact that we regard each hypersimplicial decorated ordered set partition up to cyclic rotation of blocks.
Let $\Gamma_{2,n}$ be the graph of alcoved triangulation of the hypersimplex $\Delta_{2,n}$, see Definition 2.4. We associate a hypersimplicial decorated ordered set partition of type $(2,n)$ with winding number 1 to each edge in $\Gamma_{2,n}$.
Definition 6.4.
Let $A,A^{\prime}\subseteq\Delta_{2,n}$ be two alcoves such that $A$ and $A^{\prime}$ share a common facet. Then $A\cap A^{\prime}$ is of the form $y_{j}-y_{i}=1$ for some $i\not\equiv j\pm 1\bmod n$. We associate the hypersimplicial decorated ordered set partition $\{\{i,\dots,j-1\},\{j,\dots,i-1\}\}$ to the edge $(A,A^{\prime})$ in the graph $\Gamma_{2,n}$.
Definition 6.5.
Two hypersimplicial decorated ordered set partitions of type $(2,n)$ and winding number 1 are adjacent if they can be associated to facets of the same alcove.
Example 6.6.
[…]
To relate the formula in Theorem 5.4 to hypersimplicial decorated ordered set partitions, we define a map $\psi$ from a set of $d$ adjacent hypersimplicial decorated ordered set partitions of winding number 1 to a single hypersimplicial decorated ordered set partition of winding number $d$.
Definition 6.7.
For two hypersimplicial decorated ordered set partitions $\{A,B\},\{A^{\prime},B^{\prime}\}$ of type $(2,n)$ with nonzero winding numbers, we define $\psi(\{A,B\},\{A^{\prime},B^{\prime}\})$ to be a new hypersimplicial decorated ordered set partition $\{A^{\prime\prime},B^{\prime\prime}\}$ of type $(2,n)$ such that $A^{\prime\prime}=A\triangle A^{\prime}$ is the symmetric difference between $A$ and $A^{\prime}$ and $B^{\prime\prime}=[n]\setminus A^{\prime\prime}$ is the complement of $A^{\prime\prime}$.
For a collection of $d$ adjacent hypersimplicial decorated ordered set partitions $\{A_{1},B_{1}\},\dots,\{A_{d},B_{d}\}$ of type $(2,n)$ and winding number 1, we define $\psi(\{A_{1},B_{1}\},\dots,\{A_{d},B_{d}\})$ recursively to be
$\psi(\psi(\{A_{1},B_{1}\},\dots,\{A_{d-1},B_{d-1}\}),\{A_{d},B_{d}\}).$
Example 6.8.
[…]
Example 6.9.
Consider $\{\{1,2,3\},\{4,5,6\}\}$ and $\{\{2,3,4\},\{5,6,1\}\}$ of type $(2,6)$ and winding number 1, then $\psi(\{\{1,2,3\},\{4,5,6\}\},\{\{2,3,4\},\{5,6,1\}\})=\{\{1,4\},\{2,3,5,6\}\}$.
Example 6.10.
Consider $\{\{1,2,3\},\{4,5,6\}\}$, $\{\{2,3,4\},\{5,6,1\}\}$, and $\{\{3,4,5\},\{6,1,2\}\}$ of type $(2,6)$ and winding number 1, we have
$\psi(\{\{1,2,3\},\{4,5,6\}\},\{\{2,3,4\},\{5,6,1\}\},\{\{3,4,5\},\{6,1,2\}\})$
$=\psi(\{\{1,4\},\{2,3,5,6\}\},\{\{3,4,5\},\{6,1,2\}\})$
=$=\{\{1,3,5\},\{2,4,6\}\}.$
Conjecture 6.11.
For any $n$ and for any alcove $A_{0}$ in $\Delta_{2,n}$, let $\mathcal{P}_{A_{0}}=(V,E)$ be the breadth-first search order of $\Gamma_{2,n}$ with root $A_{0}$ (see Definition 5.1). For $v\in V$, let $\operatorname{cover}(v)=\#\{u\in V\mid u\prec\cdot~{}v\text{ in }\mathcal{P}_{A_{0}}\}$ be the number of elements $v$ covers. The map $\psi$ is a bijection from the set $\{v\in V\mid\operatorname{cover}(v)=d\}$ and the set of hypersimplicial decorated ordered set partitions of type $(2,n)$ with winding number $d$.
$\{3,5\}_{1},\{1,2,4\}_{1}$$\{2,4\}_{1},\{1,3,5\}_{1}$$\{1,3\}_{1},\{2,4,5\}_{1}$$\{2,5\}_{1},\{1,3,4\}_{1}$$\{1,4\}_{1},\{2,3,5\}_{1}$$\{3,4\}_{1},\{1,2,5\}_{1}$$\{2,3\}_{1},\{1,4,5\}_{1}$$\{1,2\}_{1},\{3,4,5\}_{1}$$\{1,5\}_{1},\{2,3,4\}_{1}$$\{4,5\}_{1},\{1,2,3\}_{1}$$\{1,2,3,4,5\}_{2}$
Figure 4. This is the graph of alcoved triangulation of the hypersimplex $\Delta_{2,5}$. Our root alcove $A_{0}$ is at the center. The arrows indicate cover relations in $\mathcal{P}_{A_{0}}$, pointing in increasing directions. The orange arrows are facets representing the cover relations of the alcove labeled by $\{1,3\}_{1},\{2,4,5\}_{1}$.
Example 6.12.
In Figure 4, we choose $A_{0}$ to be the simplex in the center of the graph $\Gamma_{2,5}$ and we label it by the unique hypersimplicial decorated ordered set partition of type $(2,5)$ with winding number 0, which is $\{1,2,3,4,5\}_{2}$. The arrows are cover relations, and they point in increasing directions in $\mathcal{P}_{A_{0}}$.
The alcove labeled by $\{1,3\}_{1},\{2,4,5\}_{1}$ covers two alcoves in the poset $\mathcal{P}_{A_{0}}$, through facets colored by orange labeled by $\{1,2\}_{1},\{3,4,5\}_{1}$ and $\{2,3\}_{1},\{1,4,5\}_{1}$. This agrees with previous Example 6.8.
[…]